# TransVIP: Speech to Speech Translation System with Voice and Isochrony Preservation

### v1

@@ B.2 Ablation on the Choice of Codec
Additionally, we find that when trained with DAC, the joint translation model fails to generate reasonable output even when the training loss appears to converge. To investigate this, we plot the validation loss curve of the joint translation model in Figure 3. It shows that SpeechTokenizer has the lowest loss, followed by SASC, and then DAC. Furthermore, we observed that SpeechTokenizer can generate reasonable output with minimal training steps. We attribute this finding to two reasons: 1) SpeechTokenizer contains most semantic information in its first-layer codec. Although SASC also undergoes semantic distillation, it still contains acoustic information in the first layer, trading for better speech reconstruction quality. 2) The enhancement of RVQ in DAC, which is also utilized in SASC, leads to a evenly utilization rate across the codebook. This could pose a challenge for language modeling. However, our SASC method effectively balances between the quality of reconstruction and the affinity of the language model.
@@
B.3 Ablation Study on the NAR Acoustic Model
The textless model is trained on a combination of datasets including Librilight, VoxPopuli French subset, SeamlessAlign, and Common Voice English and French subsets. As Librilight and VoxPopuli do not provide text transcriptions, we substitute the LibriHeavy dataset in the model trained with text input. Additionally, in the textless model, long audio segments are randomly clipped, whereas in the model with text input, clips strictly adhere to timestamps provided in the metadata to align with text transcriptions. We conduct the comparison on the Fr-En uni-directional setup to mitigate the impact of the absence of a large-scale French dataset on the model with text input. Despite the apparent discrepancy in datasets, textless modeling demonstrates superior performance, underscoring one of its key data accessibility advantages. All other hyperparameters, including training steps, remain consistent across both models.
NAR Model Datasets SIM BLEU Nat.
NAR w/o text LL,VP,SA,CV 0.320 32.60 3.19
   - LBS 0.309 32.30 3.17
NAR w/ text LH,SA,CV 0.307 31.52 3.10
   - LBS 0.298 31.03 3.09
Appendix C Pseudo code for Layer Beam Search
The pseudo code, which is written in the Pytorch style, is shown in algorithm 1.
@@ 27            $inputs\leftarrow next\_inputs$
Appendix D Model size
We report in Table 6 the detailed decomposition of modules sizes of TransVIP and baseline models.
Table 6: Parameters of the building components. Only the modules that is required in inference stage is calculated.
@@ Appendix D Model size
Appendix E Data Preprocessing and Pseudo-labeling
As mentioned in Section 3.1.1, text labels play a crucial role in joint translation model training. However, text labels are absent in some end-to-end S2ST datasets such as SeamlessAlign and SpeechMatrix. To utilize these datasets, pseudo text labels are required. We employ the following procedure to preprocess such datasets:
@@ Appendix E Data Preprocessing and Pseudo-labeling
Appendix F Subjective Evaluation
@@
Appendix G Broader Impact
TransVIP was developed to improve cross-lingual communication and assist with the automation of video dubbing workflows. However, it carries potential risks, including the misuse of the model for impersonating specific speakers by inputting targeted text and speaker prompts. Additionally, as a generative model, there is a possibility it may produce toxic or biased outputs, although such occurrences were not observed in our test samples. It is imperative to monitor for any misuse or harmful outputs rigorously. We must also provide users with the ability to report such incidents, ensuring responsible usage and continuous improvement of the system.

### v2

@@ Appendix C Additional Ablation on the Choice of Codec
Additionally, we find that when trained with DAC, the joint translation model fails to generate reasonable output even when the training loss appears to converge. To investigate this, we plot the validation loss curve of the joint translation model in Figure 3. It shows that SpeechTokenizer has the lowest loss, followed by SASC, and then DAC. Furthermore, we observed that SpeechTokenizer can generate reasonable output with minimal training steps. We attribute this finding to two reasons: 1) SpeechTokenizer contains most semantic information in its first-layer codec. Although SASC also undergoes semantic distillation, it still contains acoustic information in the first layer, trading for better speech reconstruction quality. 2) The enhancement of RVQ in DAC, which is also utilized in SASC, leads to an even utilization rate across the codebook. This could pose a challenge for language modeling. However, our SASC method effectively balances the quality of reconstruction and the affinity of the language model.
@@
Appendix D Pseudo code for Layer Beam Search
The pseudo code in the Pytorch style is shown in algorithm 1.
@@ 27            $inputs\leftarrow next\_inputs$
Appendix E Model size
We report in Table 7 the detailed decomposition of module sizes of TransVIP and baseline models.
Table 7: Parameters of the building components. Only the modules that are required in the inference stage are calculated.
@@ Appendix E Model size
Appendix F Data Preprocessing and Pseudo-labeling
As mentioned in section 3.2, text labels play a crucial role in joint translation model training. However, text labels are absent in some end-to-end S2ST datasets such as SeamlessAlign and SpeechMatrix. To utilize these datasets, pseudo text labels are required. We employ the following procedure to preprocess such datasets:
@@ Appendix F Data Preprocessing and Pseudo-labeling
Appendix G Subjective Evaluation